# envname-P envname#1

### v2

@@ 5.2 Math Reasoning
Settings. To evaluate MISA on math reasoning tasks, we tested LLaMA3-8B and Qwen2.5-7B on four arithmetic benchmarks: GSM8K [cobbe2021gsm8k], SVAMP [patel2021nlp], AQUA [ling2017program], and MAWPS [koncel2016mawps], following [hu2023llm] for dataset settings. The models were fine-tuned on MATH10K [hu2023llm], which combines training data from AQUA, GSM8K, and MAWPS, incorporating LM-generated chain-of-thought [wei2022chain] steps.
@@ 5.3 Instruction Tuning
To demonstrate MISA’s efficiency in instruction-following fine-tuning, we fine-tuned TinyLLaMA[zhang2024tinyllama], LLaMA2-7B[touvron2023llama2openfoundation] and Mistral-7B[jiang2023mistral] on the Alpaca GPT-4 dataset, which consists of 52K samples generated by GPT-4 based on inputs from the Alpaca dataset [alpaca]. Model performance was evaluated on MMLU[hendrycks2020measuring], MMLU-pro[wang2024mmlu] and MT-Bench [zheng2023judging], As shown in Table 5, MISA outperforms all baselines in most evaluation benchmarks, demonstrating its robustness and efficiency for instruction fine-tuning across different models. In addition, we compared the validation loss of MISA with other BCD methods, BAdam and LISA. As shown in Figure 3, the x-axis represents training time in minutes. MISA and LISA have similar total training time, while BAdam is faster than both. Despite this, MISA exhibits much better convergence compared to LISA and BAdam.
Table 5: Comparison of different methods on MMLU, MMLU-pro and MT-Bench. Results of LISA, GaLore, and LoRA in MMLU and MT-Bench are cited from LISA[pan2024lisa], while others are reproduced by us. MT-Bench evaluation was conducted using GPT-4. Memory was tested with batch size 2.
@@ 5.4 Pre-training
We evaluate MISA for pre-training LLaMA models [touvron2023llama2openfoundation] on the C4 dataset [raffel2020exploring]. We pre-trained the 130M and 350M variant of LLaMA2 [lialin2023relorahighranktraininglowrank]. Unlike fine-tuning, where the embedding and output layers remain frozen, we trained them using the standard Adam optimizer. The number of training tokens is 2.75B. We report results for Adam, GaLore, and MISA. Table 6 reports the validation perplexity after 52K training steps. With a high-rank subspace (i.e., $r=256$, $\delta=25\%$), MISA outperforms GaLore and Adam on the 130M model. While on the 350M model, MISA outperforms GaLore and is slightly behind Adam. As shown in Fig. 4, MISA converges better than GaLore, and is similar with Adam’s performance in the later training stage. MISA’s strong performance indicates that it can be viewed as a regularization of Adam, and many gradients are redundant during full-parameter training.
@@ B.1 Memory Efficiency When Scaling To 70B Model
Fig. 5 demonstrates that when training a 70B model, MISA is still more memory-efficient than LoRA under long sequence length settings, and the use of flash-attention[dao2023flashattention2] has minimal impact on overall memory consumption.
@@ B.2 Enhancing LoRA with MISA
A key advantage of LoRA over MISA is its ability to train separate adapters for different downstream tasks. Instead of storing full model weights, LoRA only requires storage for task-specific adapters, significantly reducing storage costs. Since MISA is orthogonal to LoRA and can be used alongside it, we explored a hybrid approach combining both methods. Given a model weight $W$ with LoRA applied, it is represented as $W^{\prime}=W+AB$, where $A$ and $B$ are treated as separate modules. We freeze the original model parameters and consider the LoRA adapters as the total set of trainable parameters $n_{\text{LoRA}}$. MISA dynamically activates LoRA adapters while ensuring that the number of active parameters at each step remains below $n_{\text{LoRA}}\times\delta$. Furthermore, we found that optimizer states are not the primary source of memory consumption for LoRA + MISA. Thus, we retain optimizer states for all LoRA adapters without clearing them. Our results show that MISA enhances full LoRA’s performance while further reducing memory usage. As shown in Fig. 6, activating no more than 30% of LoRA parameters achieves or even surpasses full LoRA’s performance while saving approximately 7.9% of memory. We also compared IST [yao2024layerwiseimportancemattersmemory] with MISA (see Table 7) and MISA outperformed IST.
@@ C.3 Impact of Inner-Loop Iteration $T$
As shown in Table 9, varying $T$ from 5 to 500 does not significantly affect the convergence of validation loss or the MMLU performance. This observation aligns with the findings in BAdam [luo2024badam] and LISA [pan2024lisa], confirming that MISA’s convergence is robust to the choice of TT—a result supported by our theoretical analysis in Theorem 1, where MISA is proven to converge for any number of inner-loop updates $T$. Notably, when $T\geq 200$, we observe a slight degradation in MMLU score (e.g., 46.01 at $T=200$ and. 46.27 at $T=50$), which may stem from over-fitting to the sampled module during extended inner-loop updates. Thus, we recommend setting T=30T=305050 in practice to balance computational efficiency and performance.

### v3

@@ 5.2 Math Reasoning
Settings. To evaluate MISA on math reasoning tasks, we tested LLaMA3-8B and Qwen2.5-7B on four arithmetic benchmarks: GSM8K [8], SVAMP [42], AQUA [34], and MAWPS [28], following [25] for dataset settings. The models were fine-tuned on MATH10K [25], which combines training data from AQUA, GSM8K, and MAWPS, incorporating LM-generated chain-of-thought [57] steps.
@@ 5.3 Instruction Tuning
To demonstrate MISA’s efficiency in instruction-following fine-tuning, we fine-tuned TinyLLaMA[66], LLaMA2-7B[53] and Mistral-7B[26] on the Alpaca GPT-4 dataset, which consists of 52K samples generated by GPT-4 based on inputs from the Alpaca dataset [52]. Model performance was evaluated on MMLU[21], MMLU-pro[56] and MT-Bench [71], As shown in Table 5, MISA outperforms all baselines in most evaluation benchmarks, demonstrating its robustness and efficiency for instruction fine-tuning across different models. In addition, we compared the validation loss of MISA with other BCD methods, BAdam and LISA. As shown in Figure 3, the x-axis represents training time in minutes. MISA and LISA have similar total training time, while BAdam is faster than both. Despite this, MISA exhibits much better convergence compared to LISA and BAdam.
Table 5: Comparison of different methods on MMLU, MMLU-pro and MT-Bench. Results of LISA, GaLore, and LoRA in MMLU and MT-Bench are cited from LISA[41], while others are reproduced by us. MT-Bench evaluation was conducted using GPT-4. Memory was tested with batch size 2.
@@ 5.4 Pre-training
We evaluate MISA for pre-training LLaMA models [53] on the C4 dataset [46]. We pre-trained the 130M and 350M variant of LLaMA2 [33]. Unlike fine-tuning, where the embedding and output layers remain frozen, we trained them using the standard Adam optimizer. The number of training tokens is 2.75B. We report results for Adam, GaLore, and MISA. Table 6 reports the validation perplexity after 52K training steps. With a high-rank subspace (i.e., $r=256$, $\delta=25\%$), MISA outperforms GaLore and Adam on the 130M model. While on the 350M model, MISA outperforms GaLore and is slightly behind Adam. As shown in Fig. 4, MISA converges better than GaLore, and is similar with Adam’s performance in the later training stage. MISA’s strong performance indicates that it can be viewed as a regularization of Adam, and many gradients are redundant during full-parameter training.
@@ B.1 Memory Efficiency When Scaling To 70B Model
Fig. 5 demonstrates that when training a 70B model, MISA is still more memory-efficient than LoRA under long sequence length settings, and the use of flash-attention[9] has minimal impact on overall memory consumption.
@@ B.2 Enhancing LoRA with MISA
A key advantage of LoRA over MISA is its ability to train separate adapters for different downstream tasks. Instead of storing full model weights, LoRA only requires storage for task-specific adapters, significantly reducing storage costs. Since MISA is orthogonal to LoRA and can be used alongside it, we explored a hybrid approach combining both methods. Given a model weight $W$ with LoRA applied, it is represented as $W^{\prime}=W+AB$, where $A$ and $B$ are treated as separate modules. We freeze the original model parameters and consider the LoRA adapters as the total set of trainable parameters $n_{\text{LoRA}}$. MISA dynamically activates LoRA adapters while ensuring that the number of active parameters at each step remains below $n_{\text{LoRA}}\times\delta$. Furthermore, we found that optimizer states are not the primary source of memory consumption for LoRA + MISA. Thus, we retain optimizer states for all LoRA adapters without clearing them. Our results show that MISA enhances full LoRA’s performance while further reducing memory usage. As shown in Fig. 6, activating no more than 30% of LoRA parameters achieves or even surpasses full LoRA’s performance while saving approximately 7.9% of memory. We also compared IST [63] with MISA (see Table 7) and MISA outperformed IST.
@@ C.3 Impact of Inner-Loop Iteration $T$
As shown in Table 9, varying $T$ from 5 to 500 does not significantly affect the convergence of validation loss or the MMLU performance. This observation aligns with the findings in BAdam [37] and LISA [41], confirming that MISA’s convergence is robust to the choice of TT—a result supported by our theoretical analysis in Theorem 1, where MISA is proven to converge for any number of inner-loop updates $T$. Notably, when $T\geq 200$, we observe a slight degradation in MMLU score (e.g., 46.01 at $T=200$ and. 46.27 at $T=50$), which may stem from over-fitting to the sampled module during extended inner-loop updates. Thus, we recommend setting T=30T=305050 in practice to balance computational efficiency and performance.